Rewrite each expression:
5+y divided by 2x(squared)+10 = ? divided by 4(xsquared + 5)
i dont really understand what to do or how to do it, please help.

Answer by stanbon(57214)

If I read you correctly you have:
[(5+y)/(2x^2+10) = ?/4(x^2+5)
Factor to get:
[(5+y)/2(x^2+5)] = ?/4(x^2+5)
Multiply both sides by (x^2+5) to get:
[(5+y)/2] = ?/4
Multiply both sides by "4" to get:
2(5+y) = ?
Somehow I doubt this is the the answer you
are looking for, but the way you posted
the problem allows for many interpretations
as to exactly what your text problem looks
like.
Maybe the procedures I have used will enable
you to correctly do the problem you have in
your text.
Cheers,
Stan H.